# Automatic dataset shift identification to support root cause analysis of AI performance drift

### v2

@@ Abstract
Shifts in data distribution can substantially harm the performance of clinical AI models. Hence, various methods have been developed to detect the presence of such shifts at deployment time. However, root causes of dataset shifts are varied, and the choice of shift mitigation strategies is highly dependent on the precise type of shift encountered at test time. As such, detecting test-time dataset shift is not sufficient: precisely identifying which type of shift has occurred is critical. In this work, we propose the first unsupervised dataset shift identification framework, effectively distinguishing between prevalence shift (caused by a change in the label distribution), covariate shift (caused by a change in input characteristics) and mixed shifts (simultaneous prevalence and covariate shifts). We discuss the importance of self-supervised encoders for detecting subtle covariate shifts and propose a novel shift detector leveraging both self-supervised encoders and task model outputs for improved shift detection. We report promising results for the proposed shift identification framework across three different imaging modalities (chest radiography, digital mammography, and retinal fundus images) on five types of real-world dataset shifts, using four large publicly available datasets.
Introduction
@@
Machine learning models are notoriously sensitive to changes in the input data distribution, a phenomenon commonly referred to as dataset shift [1, 2, 3, 4]. This is particularly problematic in clinical settings, where dataset shift is a common occurrence and may arise from various factors [5]. Changes in the frequency of disease positives over time or across geographical regions cause prevalence shift [6, 7] (also known as label shift). The use of different acquisition protocols or scanners [8, 9, 10], or a change in patient demographics [11, 12] can induce shifts in image characteristics, known as covariate shift. We illustrate examples of real-world shifts in Fig. 1. Dataset shift can dramatically affect the performance of AI and lead to clinical errors such as misdiagnosis [13, 4, 14]. It is recognised as the fundamental barrier hindering AI adoption [15, 16]. It is hence crucial to implement safeguards allowing not only effective detection of the presence of shifts, but importantly, reliable identification of the root causes. Comprehensive shift detection and identification frameworks are key for the safe deployment and continuous monitoring of AI in clinical practice.
Dataset shifts can be detected at deployment time by using statistical testing to compare the distributions of incoming test data to the distribution of the reference data (representative of the data used to validate the deployed AI model). Significant progress has been made in this field where state-of-the-art methods can detect various types of real-world shifts [17, 18, 19, 20]. Shifts between test and reference data can either be detected at the output level by comparing distributions of model outputs, or at the input level by comparing low-dimensional feature representations of input images [19] (see Methods, Dataset shift detection methods for more details). In this study, we show that different types of shifts require different shift detection approaches. On the one hand, comparing model output distributions allows for the reliable detection of shifts directly related to the downstream task, such as changes in prevalence. On the other hand, we show that for shifts orthogonal to the downstream task, such as changes in image acquisition protocols, comparing output distributions is not sufficient. For such shifts, test and reference data need to be compared at the input level using rich feature representations. We demonstrate that self-supervised neural network image encoders [21], trained without using any task-specific annotations, yield excellent low-dimensional feature representations for shift detection.
@@
Figure 1: Examples of dataset shifts in medical imaging. Reliably detecting and identifying the nature of the shift is crucial to enable the safe deployment of machine learning systems applications. In this work, we propose the first shift identification framework able to reliably detect and classify any detected shift as (i) prevalence, (ii) covariate or (iii) mixed prevalence and covariate shift.
While detecting dataset shifts is important, it is insufficient for the safe deployment of AI. Besides knowing that there is a problem, we need to be able to identify its root cause to take the necessary actions, implement preventive measures, and safeguard against harm caused by AI errors. To our knowledge, no solution yet exists to identify the root cause of dataset shifts.
@@
The precise identification of the type of shifts is critical for selecting appropriate mitigation strategies. For example, prevalence shifts can often be mitigated with lightweight output recalibration techniques [14, 22, 23, 24], but these rely on the assumption that no other types of shift are present. In contrast, covariate shifts require more advanced domain adaptation techniques or model fine-tuning [25, 26, 27, 28, 4, 29]. The difficulty is that a change in image characteristics may cause similar changes in the distribution over model outputs as a change in disease prevalence [4], and determining the cause of an observed shift can be challenging. Applying the wrong mitigation technique may, in the best case, be ineffective in resolving the shift or, in the worst case, severely harm model performance or calibration. In Appendix A, we show an example of the consequences of applying a prevalence shift correction algorithm to the wrong type of shift on model calibration, further motivating the need for dataset shift identification methods. Despite its paramount importance, automatic dataset shift identification has remained an open problem.
Figure 2: Overview of the dataset shift identification pipeline proposed in this work. We leverage both task model outputs and features from self-supervised encoders for detecting and identifying dataset shifts. Contrarily to previous works, we do not simply detect the presence of shifts but also add a second step able to identify the nature of the shift. Our method effectively separates cases of (i) prevalence shift (a change in label distribution); (ii) covariate shift (a change in image characteristics) and (iii) covariate and prevalence shift (both). Our shift identification pipeline can be divided into two stages: (A) shift detection, followed by (B) shift identification. For shift detection, we combine signals from model outputs and generic features (low-dimensional representations of images) to detect whether a shift is present in the test set (2). If a shift is detected we proceed to shift identification. Identification starts with estimating the prevalence in the test set (3), then we resample the reference set to match the estimated test set prevalence (4). We then first compare feature distributions between prevalence-adjusted reference and test set (5): if differences are no longer significant after adjusting the prevalence, the shift is attributed to prevalence shift. Conversely, if differences persist after adjusting the prevalence, then covariate shift is necessarily present. In this case, we compare model output distributions to determine whether prevalence shift is also present (6). Precisely, if there were significant differences in model output distributions before adjusting the prevalence, but this shift disappears after adjusting the prevalence, we know that prevalence shift is also responsible for the observed shift, in this case we conclude that the observed shift is a case of mixed shift (prevalence + covariate shifts). Else, we conclude that the shift is attributed to covariate shift only.
@@
In this work, we address this issue by proposing the first dataset shift identification framework capable of identifying the root cause of the underlying shift, effectively separating (i) prevalence shift, (ii) covariate shift and (iii) mixed shift (both prevalence and covariate shifts). Our shift identification framework consists of two stages and is summarised in Fig. 2. First, the ‘shift detection’ module identifies whether a shift is present. Secondly, given that a shift has been detected, our ‘shift identification’ module characterises the type of shift. An in-depth evaluation across three different clinical applications (chest radiography, digital mammography, and retinal fundus images) on five types of real-world dataset shifts demonstrates that our framework accurately distinguishes between prevalence shifts, covariate shifts, and mixed shifts across various scenarios. We conclude that the proposed method may play an important role towards safe deployment and clinical adoption of medical imaging AI.
Results
Datasets and shift generation
We evaluate our methods on four different datasets covering three different imaging modalities: (i) chest radiography using PadChest [30] and RSNA Pneumonia Detection [31] datasets; (ii) mammography using the EMBED [32] dataset, and (iii) fundus images using the Kaggle Diabetic Retinopathy Detection [33], Kaggle Aptos Blindness Detection [34] and Messidor-v2 [35] datasets. To study prevalence shift detection, we associate each dataset with a downstream task. For chest radiography datasets we focus on pneumonia detection (binary), for mammography on breast density assessment (4 classes), and for retinal images on diabetic retinopathy assessment (binary). More details can be found in Methods, Datasets.
For every dataset, we study different types of shifts at different levels of intensity. We simulate various levels of prevalence shift by resampling the test set according to specific label distributions. Similarly, we simulate different types of covariate shifts by resampling the datasets according to specific scanner, acquisition site or gender distributions. We then measure shift detection (resp. identification) accuracy by repeated sampling of shifted test sets. Specifically, for each bootstrap sample, a test set is sampled according to the defined shift generating process. We then measure how many times the shift is detected (resp. the correct shift was identified) out of all 200 bootstrap samples. Technical details on shift generation processes, experimental setup, and implementation of shift detection algorithms can be found in Methods, Shift generation details and Implementation details. This extensive experimental setup not only allows us to test the effectiveness of our shift identification framework but also to evaluate the sensitivity of existing dataset shift detection methods on a large variety of realistic shifts across modalities that have not been studied before in this context.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 3: Prevalence shift: shift detectors comparison. We report the shift detection rate over 200 bootstrap samples. Results show that output-based detection is best for detecting prevalence shifts. For each combination of test set size and type of shift, the detector with the highest detection rate, as well as all detectors not significantly different from the best, are denoted with an asterisk * .
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Refer to caption(e)
Figure 4: Covariate shifts: shift detectors comparison. We studied two sub-types of covariate shifts: subpopulation shift (a-b) and acquisition shift (c-e). We report the shift detection rate over 200 bootstrap samples. Results show that feature-based detection is best for this type of shift. For each combination of test set size and type of shift, the detector with the highest detection rate, as well as all detectors not significantly different from the best, are denoted with an asterisk * .
Different shifts require different detectors
@@ Results
We first investigate which types of shifts are successfully detected by prominent dataset shift detection methods. We compare two families of shift detectors: model output-based and feature-based detectors leveraging representations from pretrained neural networks. We additionally propose and test a dual approach that combines both complementary approaches for improved shift detection (‘Duo’). A detailed description of the shift detection tests used in this study can be found in Methods, Dataset shift detection methods.
@@
For feature-based shift detection, any pretrained network could be used as feature extractor. A perhaps obvious choice is to simply use the encoder from the task-specific classification model. However, this may not be the best choice as learned features will be heavily skewed towards encoding characteristics specifically relevant to that task as opposed to encoding more generic image representations sensitive to data distribution changes [36]. Hence, we here explore the potential of self-supervised (SSL) image encoders for shift detection. Indeed, encoders trained in a self-supervised manner, i.e. without any labels, learn fine-grained representations effectively summarising all the information encoded in a given image, resulting in ideal candidates for generic shift detection. We compare the performance of feature-based shift detection for five different encoders (using the features obtained from the last layer of the neural network encoder). We compare: (i) ‘Random’ a neural network (ResNet) encoder with random weights; (ii) ‘Supervised ImageNet’ where features are extracted using a ResNet encoder trained to perform classification on natural images from the ImageNet dataset; (iii) ‘Task model’, the ResNet encoder from the task model used to perform the downstream classification task; (iv) ‘SSL ImageNet’ a self-supervised encoder trained on ImageNet data only; (v) ‘SSL Modality Specific’ a self-supervised encoder trained on the same modality as the test datasets. Additionally, for the RETINA dataset, we include a comparison using the open-source ‘RetFound’ [37] self-supervised foundation model as a feature extractor. More details on self-supervised models can be found in Methods, Self-supervised learning.
@@
Figs. 3 and 4 show the shift detection rates for every dataset-shift combination. Across all datasets, a clear pattern appears. For prevalence shifts (Fig. 3), output-based shift detection performs significantly better than all feature-based tests. This is intuitive as the shift is directly related to the downstream prediction task. A change in prevalence should directly be reflected by a change in the distribution of task model outputs. For covariate shifts, results are very different, regardless of whether we look at acquisition or subpopulation shifts (see Fig. 4). For these shifts, most feature-based detectors perform substantially better than output-based detectors. Output-based shift detection fails to detect gender shifts on radiography datasets (less than 5% of shifts detected) and performs substantially worse than feature-based tests on acquisition shifts across all levels of shifts and datasets. For example, on EMBED with 250 test exams, the output-based shift detector only detects 25% of acquisition shifts, whereas feature-based detectors detect at least 80% of shifts (except for the random encoder). In Appendix supplementary Fig. 2 we report the false shift detection rate of shift detectors when generating test sets without shifts.
@@
Our results show that the effectiveness of feature-based dataset shift detectors is highly dependent on the choice of the encoder used to extract the features (see differences between feature-based detection rates in shades of blue in Fig. 4). The results indicate that for optimal detection of more subtle shifts (e.g. gender shifts), it is important to use an encoder trained in a self-supervised manner. The task model, for example, under-performs for gender shift detection on chest radiography datasets, and so does the encoder trained in a supervised manner on ImageNet data, particularly visible for milder shifts and smaller test set sizes. For example, both supervised encoders fail to detect the shift from 44% to 25% of females in the RSNA dataset, even with large test sizes (Fig. 4 a). Encoders trained in a self-supervised manner (SSL ImageNet and SSL Modality Specific) detect gender shifts with significantly higher sensitivity, exhibiting an average sensitivity of 85% across all gender shifts with a test set size of 250 and 100% for a test set size of 1000. Similarly, for acquisition shifts, results show that SSL encoders offer substantially better detection rates than their supervised counterparts. The SSL model trained on ImageNet data was particularly effective and, in some cases, even better than the modality-specific SSL models for PadChest and EMBED, especially for subtle shifts and in the low test data regime. On the RETINA dataset, self-supervised encoders also all outperform their supervised counterparts, with no significant differences between self-supervised encoders.
@@
Building on the previous findings, we introduce a dual detection approach combining responses from output-based and feature-based shift detectors, using self-supervised features for more robust shift detection. Results in Figs. 3 and 4 demonstrate that the proposed ‘Duo’ detector (in orange) performs best overall across shifts and datasets, regardless of the shift severity. For example, with the duo detector and test set size of 1000 samples, the detection accuracy is $>80\%$ for nearly every shift, across all datasets. On the contrary, the other detectors respectively fail either on prevalence shift (with feature-based detectors detecting less than 50% prevalence shift cases averaged over datasets) or on covariate shift (where output-based detection detects less than 5% of gender shifts for radiography and less than 25% of acquisition shifts for EMBED). We employ this dual approach for the detection module of our shift identification pipeline.
Shift identification performance

### v3

# Automatic dataset shift identification to support safe deployment of medical imaging AI

@@ Abstract
Shifts in data distribution can substantially harm the performance of clinical AI models and lead to misdiagnosis. Hence, various methods have been developed to detect the presence of such shifts at deployment time. However, the root causes of dataset shifts are diverse, and the choice of shift mitigation strategies is highly dependent on the precise type of shift encountered at test time. As such, detecting test-time dataset shift is not sufficient: precisely identifying which type of shift has occurred is critical. In this work, we propose the first unsupervised dataset shift identification framework for imaging datasets, effectively distinguishing between prevalence shift (caused by a change in the label distribution), covariate shift (caused by a change in input characteristics) and mixed shifts (simultaneous prevalence and covariate shifts). We discuss the importance of self-supervised encoders for detecting subtle covariate shifts and propose a novel shift detector leveraging both self-supervised encoders and task model outputs for improved shift detection. We show the effectiveness of the proposed shift identification framework across three different imaging modalities (chest radiography, digital mammography, and retinal fundus images) on five types of real-world dataset shifts using five large publicly available datasets. Code is publicly available at https://github.com/biomedia-mira/shift_identification.
1 Introduction
@@
Machine learning models are notoriously sensitive to changes in the input data distribution, a phenomenon commonly referred to as dataset shift [44]. This is particularly problematic in clinical settings, where dataset shift is a common occurrence and may arise from various factors [4]. Changes in the frequency of disease positives over time or across geographical regions cause prevalence shift [15, 27]. The use of different acquisition protocols or scanners [41, 35, 37], or a change in patient demographics [2, 42] can induce shifts in image characteristics, known as covariate shift. We illustrate examples of real-world shifts in Fig. 1. Dataset shift can dramatically affect AI performance of AI and may lead to clinical errors such as misdiagnosis [34, 33, 13]. It is hence crucial to implement safeguards allowing not only effective detection of the presence of shifts, but importantly, reliable identification of the root causes.
@@
Figure 1: Examples of dataset shifts in medical imaging. Reliably detecting and identifying the nature of the shift is crucial to enable the safe deployment of machine learning systems applications. In this work, we propose the first shift identification framework able to reliably detect and classify any detected shift as (i) prevalence, (ii) covariate or (iii) mixed prevalence and covariate shift, for imaging datasets.
Dataset shifts can be detected at deployment time by using statistical testing to compare the distributions of incoming test data to the distribution of the reference data (representative of the data used to validate the deployed AI model). Significant progress has been made in this field where state-of-the-art methods can detect various types of real-world shifts [24, 23, 30, 12]. Shifts between test and reference data can either be detected at the output level (by comparing distributions of model outputs), or at the input level (by comparing low-dimensional feature representations of input images) [30]. In this work, we show that different types of shifts require different shift detection approaches. On the one hand, comparing model output distributions allows for the reliable detection of shifts directly related to the downstream task, such as changes in prevalence. On the other hand, we show that, for shifts orthogonal to the downstream task, such as changes in image acquisition protocols, comparing output distributions is not sufficient. For such shifts, test and reference data need to be compared at the input level using rich feature representations. We demonstrate that self-supervised neural network image encoders, trained without using any task-specific annotations, yield excellent low-dimensional feature representations for shift detection.
@@
While detecting dataset shifts is important, it is insufficient for the safe deployment of AI. Besides knowing that there is a problem, we need to be able to identify the precise type of shift to take the necessary actions, implement preventive measures, and safeguard against harm caused by AI errors. Indeed, many domain adaptation techniques are shift-specific: applying the wrong mitigation technique may, in the best case, be ineffective in resolving the shift or, in the worst case, severely harm model performance or calibration. For example, prevalence shifts can often be mitigated with lightweight output recalibration techniques [1, 39], but these rely on the assumption that no other types of shift are present. Applying such label shift adaptation methods when the shift is actually caused by covariate shift may drastically degrade model calibration and clinical metrics. In contrast, covariate shifts require more advanced domain adaptation techniques or model finetuning [46, 21, 40]. For example, image-harmonisation techniques (e.g. [21]) or automatic correction methods (e.g. [31]) effectively mitigate effects of acquisition shifts on model performance but will fail in the case of prevalence shift. The difficulty lies in the fact that a change in image characteristics may cause similar changes in the distribution over model outputs as a change in disease prevalence [31], and determining the cause of an observed shift can be challenging. Despite its importance, automatic dataset shift identification has remained an open problem.
@@
In this work, we address this issue by proposing a dataset shift identification framework capable of identifying the root cause of the underlying shift, effectively separating (i) prevalence shift, (ii) covariate shift and (iii) mixed shift (both prevalence and covariate shifts). To the best of our knowledge, this is the first framework able to identify the type of test-time shifts in an unsupervised manner for imaging data, beyond solely detecting shifts. An in-depth evaluation across three different clinical applications (chest radiography, digital mammography, and retinal fundus images) on five types of real-world dataset shifts demonstrates that our framework accurately distinguishes between prevalence shifts, covariate shifts, and mixed shifts across various scenarios.
2 Background
2.1 Definitions of prevalence and covariate shift
Formally, let $X$ denote the input image and $Y$ denote the target (e.g. disease label). Label shift (or prevalence shift) occurs when label distribution changes across domains, i.e. $P_{ref}(Y)\neq P_{test}(Y)$, while the conditional distributions, i.e. $P_{ref}(Y|X)=P_{test}(Y|X)$ are preserved, where $P_{ref}$ and $P_{test}$ denote distributions on reference and target domains respectively. Conversely, covariate shift occurs when $P_{ref}(X)\neq P_{test}(X)$, while conditional distributions are preserved [29]. Acquisition and subpopulation shifts are cases of covariate shifts as they directly affect image appearance.
2.2 Dataset shift detection methods
Several paradigms have been proposed for dataset shift detection. The simplest method to implement consists of comparing distributions of a classifier’s outputs between the reference and test domain, proposed by Rabanser et al. [30], and referred to as Black Box Shift Detection (BBSD). In detail, softmax model outputs are collected for all samples in the reference and test sets. Then, for each class, a separate univariate Kolmogorov-Smirnov (K-S) test is run to determine if the class-wise predicted probabilities distributions differ between reference and test domain, the overall significance of the shift is then determined after applying Bonferroni correction [11] for multiple testing.
In the same study, Rabanser et al. [30] also proposed another type of shift detector where reference and test data input distribution are compared using a feature-based approach. In this test, the input sample (image) first gets projected to a smaller dimension, e.g. through a pretrained neural network encoder and the shift is then measured using the Maximum Mean Discrepancy permutation test originally proposed by Gretton et al. [16]. The Maximum Mean Discrepancy measures the distance between two distributions $P$ and $Q$ based on the distance between their mean embeddings. An unbiased estimate of the square of the MMD statistic can be computed via:
$\widehat{MMD}^{2}=\frac{1}{m^{2}-m}\sum_{i=1}^{m}\sum_{j\neq i}^{m}\kappa(z_{i},z_{j})+\frac{1}{n^{2}-n}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\kappa(z^{\prime}_{i},z^{\prime}_{j})-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\kappa(z_{i},z^{\prime}_{j}),$ (1)
where $\{z_{i}\}_{i=1}^{m}\sim P$, and $\{z^{\prime}_{i}\}_{i=1}^{n}\sim Q$ and $\kappa$ is a kernel over the embedding space. The p-value can then be obtained using a permutation test. In Rabanser et al. [30], they proposed to use the RBF kernel $\kappa(z,\tilde{z})=e^{-\frac{1}{2\sigma}||z-\tilde{z}||^{2}}$, setting $\sigma$ as the median distance between all samples. An alternative way would be to explicitly learn the kernel, also known as ‘deep kernels’ [25]. The disadvantage of this approach is that the kernel needs to be learned for every single test set for which one wishes to run shift detection. In this work, we use the RBF kernel.
In the medical imaging domain, all these shift detectors were benchmarked by Koch et al. [23] for their ability to detect a large variety of real-world shifts in the context of diabetic retinopathy grading models. Both the domain classifier and the BBSD approach were shown to successfully detect the tested shifts, their sensitivity depending on the number of test samples, whereas MMD-based tests were found to be less accurate at detecting shifts compared to BBSDs and domain classifiers. Given these results, and for computational efficiency, we herein focus on shift detection methods that do not require the training of additional models at test-time.
3 Methods and experimental setup
3.1 Dataset shift identification pipeline
Figure 2: Overview of the proposed dataset shift identification pipeline. We leverage both task model outputs and features from self-supervised encoders for detecting and identifying dataset shifts. Contrarily to previous works, we do not simply detect the presence of shifts but also add a second step able to identify the nature of the shift. Our method effectively separates cases of (i) prevalence shift (a change in label distribution), (ii) covariate shift (a change in image characteristics) and (iii) covariate and prevalence shift (both).
Here, we propose a framework for identifying whether dataset shift is caused by prevalence shift, by covariate shift or by a mix of both. The approach is divided into two stages, as depicted in Fig. 2:
  1. 1.
    Standard dataset shift detection to separate the ‘shift’ from the ‘no shift’ cases (Fig. 2, A). For this step, we use a dual detection approach, combining signals from task model outputs and features from self-supervised (SSL) encoders to detect shifts. In detail, we first independently run the BBSD and the MMD shift detection tests; this yields $C$ p-values for the BBSD test (one per class) and one p-value for the MMD permutation test. We then apply Bonferroni correction on the $C+1$ p-values to get the overall significance.
  2. 2.
    If a shift is detected in the first step, we then proceed to shift identification. This process starts with estimating the prevalence in the test set (Fig. 2, B.3). For this, we can leverage the prevalence shift adaptation literature, where various methods have been proposed to estimate the density ratio:
    $w:=\frac{P_{ref}(Y)}{P_{test}(Y)}$ (2)
    with $w\in\mathbb{R}^{C}$ [1, 32, 39]. Here, we use the state-of-the-art ‘class probability matching with calibrated network’ (CPMCN) method by Wen et al. [39] to estimate this ratio and the test set prevalence. In CPMCN, the density ratio is estimated by:
    $\hat{w}:=\arg\min_{w\in\mathbb{R}^{C}}\sum_{i=1}^{C}\left|\hat{P}_{ref}(Y=i)-\frac{1}{m}\sum_{x\in D_{t}}\frac{\hat{p}(i|x)}{\sum_{j}^{C}w_{j}\hat{p}(j|x)}\right|^{2}$ (3)
    $\hat{P}_{test}(Y=i)=\hat{w}_{i}\cdot\hat{P}_{ref}(Y=i),\forall i\in C$ (4)
    We follow this method to estimate the label distribution in the test set in our shift identification module. Next, we resample the reference set to match this estimated prevalence (Fig. 2, B.4). We then first compare feature distributions between prevalence-adjusted reference and test set (Fig. 2, B.5): if differences are no longer significant after adjusting the prevalence, the shift is attributed to prevalence shift. Conversely, if differences persist after adjusting the prevalence, then covariate shift is necessarily present. In this case, we compare model output distributions with BBSD to determine whether prevalence shift is also present (Fig. 2, B.6). Precisely, if there were significant differences in model output distributions before adjusting the prevalence, but this shift disappears after adjusting the prevalence, we know that prevalence shift is also responsible for the observed shift, in this case we conclude that the observed shift is a case of mixed shift (prevalence + covariate shift). Else, we conclude that the shift is attributed to covariate shift only.
3.2 Datasets
We evaluate our methods on four different datasets covering three different imaging modalities: (i) chest radiography, (ii) mammography, and (iii) fundus images.
For chest radiography, we use two public datasets for our analysis: the RSNA Pneumonia dataset [36], a subset of the NIH Chest-Xray8 dataset [38] manually relabelled by expert radiologists for presence of pneumonia-like opacities. We use the original metadata from the NIH Chest-Xray8 dataset [38] to retrieve patient gender. We also use PadChest [3] a larger dataset containing multiple disease labels extracted from radiology reports. We here focus on the pneumonia label. Importantly, this dataset contains important metadata such as scanner information or the gender of the patient, allowing us to generate a wide range of shifts. For mammography, we use the EMBED dataset [20] a large mammogram dataset collected in the US, on six different scanners. Finally, for fundus imaging, we create RETINA a multi-domain dataset by combining three different public datasets: the Kaggle Diabetic Retinopathy Detection [10] dataset, the Kaggle Aptos Blindness Detection dataset [22] and the Messidor-v2 dataset [8]. These datasets cover different regions of the world (India, France, US) but also with varying image acquisition devices: images from the Messidor-v2 are high-quality images, while many images in the Kaggle datasets are of lower quality (including phone pictures). Creating this multi-centre dataset allows us to simulate various domain shifts by varying the proportion of data from each ‘site’ (original dataset source). The task of interest for fundus images here is binary diabetic retinopathy (DR) classification for fundus images, where we classify images between referable DR (grades 2,3,4) and healthy/non-referable DR (grades 0,1).
3.3 Shift generation details
For every dataset, we study different types of shifts at different levels of intensity. To study prevalence shift detection, we associate each dataset with a downstream task. For chest radiography datasets we focus on pneumonia detection, for mammography on breast density assessment (4 classes), and for retinal images on binary diabetic retinopathy classification. We simulate various levels of prevalence shift by resampling the test set according to specific label distributions. Then, we study various types of covariate shifts. For PadChest, we study gender shifts by varying the proportion of female patients in the test set. Moreover, PadChest contains scans acquired with two scanners, ‘Phillips’ (40%) and ‘Imaging’ (60%). This allows the simulation of different levels of acquisition shift by varying the proportion of Phillips scans in the test set. Similarly, for EMBED [20], we study acquisition shift by varying the distribution of scanners in the test set. This dataset offers a complementary view to PadChest, with a multi-class task of interest and providing even more flexibility for simulating diverse acquisition shifts (six scanners). Note that, in EMBED, each exam comprises four mammograms (left/right breasts and MLO/CC views). We excluded all exams that did not contain exactly four images, kept exactly one exam per patient, and ensured that test set sampling was done at the exam level. Finally, for the RETINA dataset, we simulate covariate shifts by varying the proportion of samples coming from each underlying dataset (Aptos, Kaggle DR and Messidor).
3.4 Implementation details
To evaluate the shift identification accuracy of the proposed framework, we follow standard evaluation practices from the dataset shift detection literature. Specifically, we repeat the following process 200 times: (i) sample a subset of size $N_{test}$ from the test split according to the shift of interest and sample the reference set from the validation split, (ii) run the shift identification test, and (iii) record whether the shift is correctly identified. The final shift identification accuracy is computed as the proportion of times the shift is correctly identified out of all the bootstrap samples. For both chest radiography datasets, we used $N_{ref}=2,000$ images and $N_{test}\in\{100,250,500,1000\}$ images. For the RETINA dataset, we used $N_{ref}=1000$ images and $N_{test}\in\{100,250,1000\}$ images. The case of EMBED is slightly different as we need to ensure that images belonging to the same exam are always sampled simultaneously. Hence, for this dataset sampling of all the reference sets, shifted test sets (and permutations in the MMD test [^1] Keeping the underlying structure of the data during permutation tests is important to avoid inflated type I error [7].) is done at the exam level and not at the image level. We assumed $N_{exams,ref}=1000$ exams (i.e $4000$ images) and $N_{exams,test}\in\{50,100,250\}$.
All task models and encoders used in this study are ResNet-50 [18]. For each dataset, we use the task model for the BBSD test. For the MMD detection test, we first extract the embeddings from the last layer of the encoder, then project them onto the 32-first principle components and then use the RBF kernel. We compare various encoders for feature extraction, in particular, encoders trained in a self-supervised manner. Self-supervised (SSL) encoders were trained using the SimCLR [5] objective. Additionally, for the RETINA dataset, we compare with RetFound [45] a publicly available self-supervised foundation model trained with the Masked Auto Encoder [17] objective on a large set of retinal images, based on a vision transformer architecture [9].
4 Results
4.1 Different shifts require different detectors
@@
Prior to diving into shift identification, we first investigate which types of shifts are successfully detected by prominent dataset shift detection methods. We compare two families of shift detectors: model output-based (BBSD) and feature-based detectors (MMD). We additionally test a dual approach that combines both approaches for improved shift detection (‘Duo’, as described in Section 3.1).
Figure 3: Prevalence shift: shift detectors comparison. We report the shift detection rate over 200 bootstrap samples. Results show that output-based detection is best for detecting prevalence shifts. For each combination of test set size and type of shift, the detector with the highest detection rate, as well as all detectors not significantly different from the best, are denoted with an asterisk * (where significance is measured by Fisher’s exact test, at level .05, with Bonferonni correction applied for multiple testing).
Figure 4: Covariate shifts: shift detectors comparison. We studied two sub-types of covariate shifts: subpopulation shift (top two rows) and acquisition shift (bottom three rows). We report the shift detection rate over 200 bootstrap samples. Results show that feature-based detection is best for this type of shift. For each combination of test set size and type of shift, the detector with the highest detection rate, as well as all detectors not significantly different from the best, are denoted with an asterisk * (where significance is measured by Fisher’s exact test, at level .05, with Bonferonni correction applied for multiple testing).
@@
For feature-based shift detection, any pretrained network could be used as feature extractor. A perhaps obvious choice is to simply use the encoder from the task-specific classification model. However, this may not be the best choice as learned features will be heavily skewed towards encoding characteristics specifically relevant to that task as opposed to encoding more generic image representations sensitive to data distribution changes [43]. Hence, we here explore the potential of SSL image encoders for shift detection. Indeed, encoders trained in a self-supervised manner, i.e. without any labels, learn fine-grained representations effectively summarising all the information encoded in a given image, resulting in ideal candidates for generic shift detection. We compare the performance of feature-based shift detection for five different encoders (using the features obtained from the last layer of the neural network encoder). We compare: (i) Random a ResNet-50 encoder with random weights; (ii) Supervised ImageNet where features are extracted using a ResNet encoder trained to perform classification on natural images from the ImageNet dataset; (iii) Task model, the ResNet encoder from the task model used to perform the downstream classification task; (iv) SSL ImageNet a self-supervised encoder trained on ImageNet data only; (v) SSL Modality Specific a self-supervised encoder trained on the same modality as the test datasets. Additionally, for the RETINA dataset, we include a comparison using the RetFound foundation model as a feature extractor [45].
@@
Figs. 3 and 4 show the shift detection rates for every dataset-shift combination. Across all datasets, a clear pattern appears. For prevalence shifts (Fig. 3), output-based shift detection performs significantly better than all feature-based tests. This is intuitive as the shift is directly related to the downstream prediction task. A change in prevalence should directly be reflected by a change in the distribution of task model outputs. For covariate shifts, results are very different, regardless of whether we look at acquisition or subpopulation shifts (see Fig. 4). For these shifts, most feature-based detectors perform substantially better than output-based detectors. Output-based shift detection fails to detect gender shifts on radiography datasets (less than 5% of shifts detected) and performs substantially worse than feature-based tests on acquisition shifts across all levels of shifts and datasets. For example, on EMBED with 250 test exams, the output-based shift detector only detects 25% of acquisition shifts, whereas feature-based detectors detect at least 80% of shifts (except for the random encoder).
@@
Our results show that the effectiveness of feature-based dataset shift detectors is highly dependent on the choice of the encoder used to extract the features (see differences between feature-based detection rates in shades of blue in Fig. 4). The results indicate that for optimal detection of more subtle shifts (e.g. gender shifts), it is important to use an encoder trained in a self-supervised manner. The task model, for example, under-performs for gender shift detection on chest radiography datasets, and so does the encoder trained in a supervised manner on ImageNet data, particularly visible for milder shifts and smaller test set sizes. For example, both supervised encoders fail to detect the shift from 44% to 25% of females in the RSNA dataset, even with large test sizes (Fig. 4 top row). Encoders trained in a self-supervised manner (SSL ImageNet and SSL Modality Specific) detect gender shifts with significantly higher sensitivity, exhibiting an average sensitivity of 85% across all gender shifts with a test set size of 250 and 100% for a test set size of 1000. Similarly, for acquisition shifts, results show that SSL encoders offer substantially better detection rates than their supervised counterparts. The SSL model trained on ImageNet data was particularly effective and, in some cases, even better than the modality-specific SSL models for PadChest and EMBED, especially for subtle shifts and in the low test data regime. On the RETINA dataset, self-supervised encoders also all outperform their supervised counterparts, with no significant differences between self-supervised encoders.
@@
Building on the previous findings, we evaluate a dual detection approach combining responses from output-based and feature-based shift detectors, using self-supervised features for more robust shift detection. Results in Figs. 3 and 4 demonstrate that the proposed ‘Duo’ detector (in orange) performs best overall across shifts and datasets, regardless of the shift severity. For example, with the duo detector and test set size of 1000 samples, the detection accuracy is $>$80% for nearly every shift, across all datasets. On the contrary, the other detectors respectively fail either on prevalence shift (with feature-based detectors detecting less than 50% prevalence shift cases averaged over datasets) or on covariate shift (where output-based detection detects less than 5% of gender shifts for radiography and less than 25% of acquisition shifts for EMBED). We employ this dual approach for the detection module of our shift identification pipeline.
For completeness, in Fig. 5 we report the false shift detection rate of shift detectors when generating test sets without shifts. We verify that the false-positive rates of the output-based, feature-based and Duo detectors stay low across datasets and detectors, we find that false detection rates hover around the expected type-I error of the statistical tests 5% for most datasets and detection methods. The only exception is EMBED, where false positive rates are a little higher than expected for the output-based (and hence Duo) detector. This is due to the fact that - for this highly imbalanced classification problem - small randomly resampled test sets may lead to small differences in the observed cumulative distribution of predictions, even in the absence of a shift in the test set generation process.
Figure 5: False detection rate: shift detectors comparison. Percentage of shift detected when resampling the test set without shift.
4.2 Shift identification performance